# Empowering Visual Artists with Tokenized Digital Assets with NFTs

### v1

@@ 6.1 User Experience Improvement
The development of all-in-one platforms that integrate various aspects of NFT operations, such as NFT creation and trading, can also enhance the user experience. These platforms should offer seamless interoperability with different blockchains and marketplaces by reducing the need for users to navigate multiple sites and services. Some features can make the NFT ecosystem more approachable, for example, automatic wallet setup, integrated educational resources, and user-friendly dashboards.
@@ 6.2 Security and Privacy Improvement
The NFT ecosystem benefits from the inherent transparency provided by blockchain technology. However, this transparency also raises potential issues. Smart contract audits by security teams can identify and prevent exploits. In addition, stronger authentication methods, such as multi-factor authentication (MFA), can significantly reduce the unauthorized access issues. For high value NFTs, cold storage, which means storing them in hardware wallets, can minimize hacking risks. Educating users about phishing scams and implementing anti-phishing measures are also crucial to protecting against attacks. Moreover, education should teach users to avoid giving away their private keys or other sensitive information. Implementing decentralized identity (DID)[165] solutions can ensure secure and verifiable identities without relying on centralized entities, preventing identity theft and fraud. Regular security updates for NFT platforms and tools are necessary to patch vulnerabilities and stay ahead of the latest security threats and best practices[35].
While blockchain transparency is a key feature, it can also compromise privacy. For artists and NFT users, maintaining privacy is crucial [10]. Zero-knowledge proofs (ZKPs)[32] can verify transactions without revealing underlying details, ensuring that transactions are transparent and verifiable while keeping the identities of the parties involved private. Confidential transaction mechanisms[201] can hide transaction amounts and other sensitive details from the public while still ensuring validity. Creating private NFTs with privacy features, achieved through encryption and selective disclosure techniques, can ensure that only authorized parties can view metadata and ownership details. Off-chain privacy solutions, where sensitive information is stored off-chain (e.g., using trusted hardware [120]) while the blockchain is used only for validation, help maintain privacy by leveraging blockchain security while keeping private data off the public ledger. Developing and promoting the use of privacy-focused wallets with features like transaction mixing [183] and address obfuscation can help users maintain their anonymity.
The continuous involvement and collaboration of artists, users, governors, and developers are essential for the ecological perfection of the NFT ecosystem. Encouraging active community participation can lead to the early identification of issues and collaborative problem-solving. Bug bounty programs can incentivize security researchers to find and report vulnerabilities. Offering education and resources about best practices for security can help users and artists protect themselves and their assets. Moreover, rules are needed to protect users and encourage innovation from regulators which can help make the NFT space more secure and trustworthy.
@@ 6.3 Remix: The Future of Next Generation NFTs
Artwork remixing involves adding user’s own ideas to an original artwork to create a new version. It’s an interesting topic that benefits both artists and the NFT market. Artists can earn ongoing royalties whenever their work is resold as an NFT remix. It also lets users modify and create new NFTs based on their own ideas, effectively turning them into new artists. Theoretically, this remixing process can potentially attract increasing numbers of active artists, as well as a variety of new creative outputs. In turn, this is likely toexpand the NFT market in ways that have yet to be considered.
Artwork remixing [106] is highly promising in the NFT world as blockchain technology ensuring artists receive payment for every resale of their artwork. This means that original artists can keep earning money and get more recognition if more of their NFTs are traded. The continuous royalty payments incentivize artists to enter and engage in the NFT market.
For users, remixing and creating new versions of NFTs allow for personal expression and creativity. By turning existing artwork into new, unique pieces, users become co-creators and contribute to the diversity of the digital art world. This open approach to art creation means that anyone with interest and creativity can join the art community, potentially bringing in a lot of new talents and innovation to the art industry.

### v2

@@ 6.1 User Experience Improvement
The development of all-in-one platforms (i.e., integrating various NFT operations, such as creation and trading) can also enhance the user experience. These platforms should offer seamless interoperability with different blockchains and marketplaces by reducing the need for users to navigate multiple services. Some features can make NFTs more approachable, for example, automatic wallet setup, integrated educational resources, and user-friendly dashboards.
@@ 6.2 Security and Privacy Improvement
NFT benefits from the inherent transparency provided by blockchain technology. However, this transparency also raises potential issues. Smart contract audits by security teams can identify and prevent exploits. In addition, stronger authentication methods, such as multi-factor authentication (MFA), can significantly reduce the unauthorized access issues. For high value NFTs, cold storage, which means storing them in hardware wallets, can minimize hacking risks. Educating users about phishing scams and implementing anti-phishing measures are also crucial to protecting against attacks. Moreover, education should teach users to avoid giving away their private keys or other sensitive information. Implementing decentralized identity (DID) [165] solutions can ensure secure and verifiable identities without relying on centralized entities, preventing identity fraud. Regular security updates for NFT platforms and tools are necessary to patch vulnerabilities and stay ahead of the latest security threats and best practices [35].
Blockchain may compromise user privacy. For artists and NFT users, maintaining privacy is crucial [10]. Zero-knowledge proofs (ZKPs) [32] can verify transactions without revealing underlying details, ensuring that transactions are transparent and verifiable while keeping the identities of the parties involved private. Confidential transaction mechanisms [201] can hide transaction amounts and other sensitive details from the public while still ensuring validity. Creating private NFTs with privacy features, achieved through encryption and selective disclosure techniques, can ensure that only authorized parties can view metadata and ownership details. Off-chain privacy solutions, where sensitive information is stored off-chain (e.g., using trusted hardware [120]) while the blockchain is used only for validation, help maintain privacy by leveraging blockchain security while keeping private data off the public ledger. Developing and promoting the use of privacy-focused wallets with features like transaction mixing [183] and address obfuscation can help users maintain their anonymity.
The continuous involvement and collaboration of artists, users, governors, and developers are essential for the ecological perfection of the NFT ecosystem. Encouraging active community participation can lead to the early identification of issues and collaborative problem-solving. Bug bounty programs can incentivize security researchers to find and report vulnerabilities. Offering education and resources about best practices for security can help artists protect their assets. Moreover, rules are needed to protect users and encourage innovation from regulators which can help make the NFT space more secure.
@@ 6.3 Remix: The Future of Next Generation NFTs
Artwork remixing involves adding user’s own ideas to an original artwork to create a new version. It is an interesting topic that benefits both artists and the NFT market. Artists can earn ongoing royalties whenever their work is resold as an NFT remix. It also lets users create new NFTs based on their own ideas, turning them into new artists. Theoretically, this remixing process can attract increasing numbers of active artists, as well as a variety of new creative outputs.
More specifically, artwork remixing [106] is promising as blockchain ensuring artists receive payment for every resale of their artwork. Original artists can keep earning money and get more recognition if more of their NFTs are traded. The continuous royalty payments incentivize artists to enter and engage in the NFT market. For users, remixing and creating new versions of NFTs allow for personal expression and creativity. By turning existing artwork into new, unique pieces, users become co-creators and contribute to the diversity of the digital art world. This open approach to art creation means that anyone with interest and creativity can join the art community, potentially bringing in a lot of new talents and innovation to the art industry.